# DNN Task Assignment in UAV Networks: A Generative AI Enhanced Multi-Agent Reinforcement Learning Approach

### v2

@@ III-F Energy Consumption Model
3) As one of the energy consumptions that must be considered, the flight energy consumption of the UAV is adopted in [30]. The flight energy consumption is approximated as the product $e_{n}^{fly}$ of the propulsion power $P^{fly}$ of the UAV and the total flying time during the task execution as follows
@@
Consequently, the energy consumption of a wingman UAV is $e_{n}=e_{n}^{comp}+e_{n}^{trans}+e_{n}^{fly}$. And then the total energy consumption of all DNN tasks executed by the UAV swarm is denoted as
@@
Path planning method based on greedy algorithm
@@ V-A Fitness Function
This paper employs a greedy algorithm to minimize the total flying distance of the UAV while jointly considering the task size of the target area to address the UAV path planning problem. Before executing the DNN task assignment, the UAV must plan an optimal flying path to reduce task completion time and energy consumption. Given that the task sizes of each target area vary, we denote the task size of the $q$-th target coordinate as $\mathbb{L}_{q}$, and let ${\upsilon}_{q}$ represent the average rate at which the UAV swarm processes tasks. Consequently, the time required to process the task size at the $q$-th target coordinate is denoted as
@@
Accordingly, the fitness function integrates the UAV flying distance and the quantity of tasks processed by the DNN model, represented as
@@
where $\vartheta$ and $\rho$ are weight coefficients utilized to balance the trade-off between the UAV flying distance and the amount of data processed by the DNN model.
@@
$\mathfrak{q}(1)=\arg\max_{\mathfrak{q}(q)\in\widetilde{\mathbb{W}}}\left(\frac{\mathbb{L}_{\mathfrak{q}}}{D_{\mathfrak{q}(q),\mathfrak{q}(0)})}\right)$
@@ V-B Path Planning Method
Specifically, the candidate target coordinates $k$ are randomly selected from the set of uninspected target coordinates, denoted as $\widetilde{\mathbb{W}}$. If $|\widetilde{\mathbb{W}}|\leq\mathrm{k}$ indicates that all uninspected target coordinates are chosen as candidate target coordinates $k$, then we define this set as $\tilde{\mathrm{g}}=\widetilde{\mathrm{W}}$. Otherwise, $\widetilde{\mathbb{W}}$ uninspected target coordinates $k$ are randomly selected from $\tilde{\mathrm{g}}\subset\widetilde{\mathbb{W}}$. For each candidate target coordinate, the fitness function value ${F}$ is calculated and the target coordinate with the smallest value $F$ is selected as the next target coordinate. A lower value of $F$ corresponds to a lower total cost. The chosen target coordinate is then added to the flying path sequence and removed from the set of candidate target coordinates. This process is repeated until all target coordinates have been inspected once, at which coordinate the UAV returns to the initial target coordinate.
@@
Figure 4: GDM-MADDPG for strategy generation of DNN task assignment.
DNN task assignment algorithm based on GDM-MADDPG
@@
VI-A Design Process of GDM-MADDPG Algorithm
DNN Task assignment decision generation differs from traditional applications of GDM, primarily because decision optimization often lacks a large, dedicated dataset for offline training. In this paper, the DNN task assignment decision process involves the GDM predicting noise distribution through iterative refinement, followed by training its reverse denoising process, specifically denoising Gaussian noise to yield optimal action decisions. This training enables the agent to execute the generated actions within the environment. The agent then fine-tunes its network parameters based on received rewards, thereby enhancing its action decision generation capabilities. This approach transforms the challenge of limited datasets into an opportunity for dynamic online learning and agent self-optimization. Building on these principles, the design process encompasses the following processes. Agent Environment Design: The types of tasks and data sizes at various target coordinates are heterogeneous and subject to dynamic changes. GDM is crafted to generate optimal assignment actions, including split point selection and task selection, based on the observations of a given agent. Action Space Design: The action space is established by applying a series of denoising steps on Gaussian noise via GDM, with the action vector dimension corresponding to the number of tasks. Reward Function Design: The objective is to derive action strategies for the assignment of DNN tasks through GDM, specifically targeting optimal solutions related to aspects such as AoI, load balancing, and energy consumption. Model Training Design: This step focuses on training the action network to enable the agent to produce specific actions based on the observations, all with the aim of maximizing the utility function.
@@ 26                        Minimize the loss function to update the parameters of the Critic network
$\mathcal{A}^{n}=\{a_{1}^{n},a_{2}^{n},\ldots a_{t}^{n}\}$. In each time slot t, the agent’s action is denoted by $a_{t}^{i}$, where $i=\{1,2,...,I\}$ represents the task chosen by the agent for execution, and $a_{t}^{n}=\{a_{t}^{1},a_{t}^{2},...,a_{t}^{i}\}$ signifies the action taken.
@@
Experiment Setup and Performance Evaluation
@@ VII-A Experiment Setup
We perform simulations in the following scenario: an airship positioned at an altitude of 10 km, accompanied by nine UAVs at an altitude of 3 km, covering an area of 12×12 km². The airship is located at the center of this area, while the leader UAV and wingman UAVs are uniformly distributed throughout it. The airship obtains task directives from the ground control center, formulates the optimal flight sequence for the designated task targets, and transmits the planned route information to the lead aircraft. The lead aircraft, along with several wingmen UAV, then executes a coordinated flight along the predetermined route. It is important to note that all UAVs fall within the coverage area of the airship and are interconnected and capable of communicating with one another. Furthermore, considering overall accuracy in detecting target amounts, the Yolov5 model outperforms the seven attention-optimized Yolov5 variants, as well as other state-of-the-art object detection models [4]. In this study, we employ Yolov5 as the task model and conduct experiments to evaluate the effectiveness of the proposed approach.
@@
3) Benchmark Algorithms
Figure 5: UAV path planning considering task size and flying distance.
@@ VII-B Performance Evaluation
This experiment simulated a UAV inspection area consisting of ten target coordinates, each defined by a distinct task size. Fig. 5a depicts the proposed the UAV path planning method. By comprehensively considering both task size and flying path, this method generates the global optimal path and target coordinate sequence, prioritizing target coordinates with larger task sizes while ensuring coverage of all target coordinates. In contrast, Fig. 5b presents a greedy algorithm which selects the best option at each step based solely on the current state, is prone to local optima. As a result, it can lead to a significant increase in flying path length, reducing the survival time of the UAV swarm and hindering its ability to conduct long-term operations under energy constraints.
Figure 6: The training procedures of different benchmark algorithms and DNN models.
Figure 7: The training procedures of different learning rate and batch size.
As illustrated in Fig. 5c, the cost associated with the optimal path finding of the UAV increases as the number of target coordinates increases. In particular, the path planning algorithm proposed in this study consistently exhibits lower costs compared to the greedy algorithm in all scenarios evaluated. Specifically, for target coordinate counts of 10, 20, 30, 40, and 50, the total cost produced by the proposed algorithm is reduced by approximately 19.3%, 21.1%, 23.0%, 25.6%, and 27.1%, respectively, compared to the greedy algorithm. This enhancement can be attributed to the expanding solution space that accompanies an increase in the number of target coordinates. As the candidate set of target coordinates grows, the proposed algorithm has greater opportunities to explore solutions that yield lower total costs, ultimately facilitating the identification of the optimal path with minimal total cost. Thus, it is evident that the algorithm developed in this study is particularly adept at addressing path planning challenges that involve a larger number of target coordinates.
@@
We perform ten independent training runs for each algorithm. As illustrated in Fig. 6a, although the proposed GDM-MADDPG converges at a slower rate than the two algorithms, GDM-MADDPG demonstrates superior performance compared to the other solutions. The total reward attained by GDM-MADDPG increases throughout the training process, eventually converging to approximately 5000 after around 600 training episodes. This enhancement is attributed to the generative capabilities of GDM, which markedly improve action sample efficiency by progressively reducing noise through multiple denoising steps. Finally, the low performance of the two algorithms further highlights the effectiveness of the proposed GDM-MADDPG method.
The convergence of GDM-MADDPG in various tasks of the DNN model is illustrated in Fig. 6b. Despite the differences in structural complexity between DNN models, the proposed method successfully accommodates these tasks and achieves convergence. The AlexNet model, characterized by fewer layers and lower computational requirements, demonstrates favorable convergence performance. In contrast, the VGG16 model, with its greater number of layers and higher computational demands, experiences increased processing delays, resulting in lower rewards and reduced convergence efficacy for our method. Although the convergence performance of our method on the YOLOv5 model is slightly inferior to that of AlexNet, YOLOv5, as a widely adopted real-time object detection algorithm, has a model complexity that lies between the two aforementioned models. Consequently, subsequent experiments employ YOLOv5 as the experimental task model to validate the superior performance of this method in terms of AOI and task completion rate.
By conducting a series of systematic experiments, we aim to identify the optimal values for two key parameters that influence the performance of the MADDPG algorithm: 1) learning rate and 2) batch size. The experimental results, illustrated in Fig. 7a, clarify the convergence behavior of the reward function under varying learning rates. A learning rate of 0.01 is excessively large, preventing convergence even after 1000 training iterations. Similarly, learning rates of 0.0001 do not facilitate convergence. In particular, the 0.001 learning rate promotes faster convergence. The suitable learning rate of 0.001 is used to assess the effects of different batch sizes on training performance. As illustrated in Fig. 7b, a batch size of 512 is recognized as the most advantageous option.
@@
This study compares the AoI of different methods in varying task sizes, as illustrated in Fig. 8a. The AoI performance of the MADDPG-based method is significantly improved when combined with the path-planning approach. Furthermore, the GDM-MADDPG method proposed in this paper, which is based on MADDPG and incorporates the advantages of GDM in handling complex data distributions and generating high-quality task assignment decisions, improves the timeliness and generalizability of task assignment decision generation. This results in the generation of the most suitable task assignment decisions for the current state, thereby reducing the transmission time during serial task execution. According to Eq. (25), this method also addresses the optimization of path planning for target coordinates and tasks, which reduces waiting times associated with tasks and improves AoI.
The task completion rates of different methods were compared in various task sizes, as shown in Fig. 8b. The method proposed in this paper outperforms other methods in optimizing system utility, indicating that optimized task path planning and task offloading decision generation achieve a joint optimization of multiple objectives. In terms of path planning, the method not only considers the optimal path but also accounts for the task-size requests at the target coordinates, enhancing the selectivity of local optima. Regarding DNN task assignment, factors such as task partitioning decisions and load balancing are incorporated. Additionally, the introduced GDM reverse denoising process improves the learning and generation capabilities of agent decisions, leading to well-generalized action decisions. Consequently, as the task size gradually increases, the MADDPG-based decision method lacks robust dynamic adaptability under multiple constraints. Its learning capability is constrained by the fixed learning mechanism of the actor-critic network, resulting in the absence of a strategy with generalization capabilities.
@@
The utility of different methods was compared under various task sizes, as shown in Fig. 8c, as task sizes increase, the need for additional task assignment work is incurred under latency constraints, directly increasing the latency of task completion. This results in a downward trend in task completion rates for all methods. The proposed method considers task size constraints during the path planning stage, and task assignment decisions can be rapidly learned and generated through the GDM method, further reducing task processing delays and allowing a greater number of tasks to be completed within the specified time. In addition, these findings indicate that the proposed method is suited for large-scale tasks.
Conclusion
In this paper, we addressed the challenge of planning joint flight paths, assigning DNN tasks, and balancing load in UAV networks as a two-stage optimization problem. In the first stage, we consider the task size of the area to be inspected and the shortest flight path as optimization constraints. We then employ a novel heuristic algorithm to optimize the UAV’s flying path, focusing on maximizing system utility. In the second stage, we introduce a novel GDM-MADDPG algorithm to determine optimal DNN task assignment decisions. Through numerical simulations, we demonstrate the superior performance of our proposed algorithm compared to existing benchmark solutions.

### v3

@@ III-F Energy Consumption Model
3) As one of the energy consumptions that must be considered, the flight energy consumption of the UAV is adopted in [30]. The flight energy consumption is approximated as the product $e_{n}^{fly}$ of the propulsion power $P^{fly}$ of the UAV and the total flight time during the task execution as follows
@@
Consequently, the energy consumption of a follow UAV is $e_{n}=e_{n}^{comp}+e_{n}^{trans}+e_{n}^{fly}$. And then the total energy consumption of all DNN tasks executed by the UAV swarm is denoted as
@@ IV Problem Formulation
V Path planning method based on greedy algorithm
@@ V-A Fitness Function
This paper employs a greedy algorithm to minimize the total flight distance of the UAV while jointly considering the task size of the target area to address the UAV path planning problem. Before executing the DNN task assignment, the UAV must plan an optimal flight path to reduce task completion time and energy consumption. Given that the task sizes of each target area vary, we denote the task size of the coordinate of the $q$th target as $\mathbb{L}_{q}$, and let ${\upsilon}_{q}$ represent the average rate at which the UAV swarm processes the tasks. Consequently, the time required to process the task size at the $q$-th target coordinate is denoted as
@@
Accordingly, the fitness function integrates the UAV flight distance and the quantity of tasks processed by the DNN model, represented as
@@
where $\vartheta$ and $\rho$ are weight coefficients utilized to balance the trade-off between the UAV flight distance and the amount of data processed by the DNN model.
@@
$\mathfrak{q}(1)=\arg\max_{\mathfrak{q}(q)\in\widetilde{\mathbb{W}}}\mathbb{L}_{\mathfrak{q}}\!\left/\!\right.\!D_{\mathfrak{q}(q),\mathfrak{q}(0)}$
@@ V-B Path Planning Method
Specifically, the candidate target coordinates $k$ are randomly selected from the set of uninspected target coordinates, denoted $\widetilde{\mathbb{W}}$. If $|\widetilde{\mathbb{W}}|\leq\mathrm{k}$ indicates that all uninspected target coordinates are chosen as candidate target coordinates $k$, then we define this set as $\tilde{\mathrm{g}}=\widetilde{\mathrm{W}}$. Otherwise, $\widetilde{\mathbb{W}}$ uninspected target coordinates $k$ are randomly selected from $\tilde{\mathrm{g}}\subset\widetilde{\mathbb{W}}$. For each candidate target coordinate, the fitness function value ${F}$ is calculated and the target coordinate with the smallest value $F$ is selected as the next target coordinate. A lower value of $F$ corresponds to a lower total cost. The chosen target coordinate is then added to the flight path sequence and removed from the set of candidate target coordinates. This process is repeated until all target coordinates have been inspected once, at which point the UAV returns to the initial target coordinate.
@@
Figure 4: GDM-MADDPG for strategy generation.
VI DNN task assignment algorithm based on GDM-MADDPG
@@
VI-A Design of GDM-MADDPG Algorithm
DNN Task assignment decision generation differs from traditional applications of GDM, primarily because decision optimization often lacks a large, dedicated dataset for offline training. In this paper, the DNN task assignment decision process involves the GDM predicting noise distribution through iterative refinement, followed by training its reverse denoising process, specifically denoising Gaussian noise to yield optimal action decisions. This approach transforms the challenge of limited datasets into an opportunity for dynamic online learning and agent self-optimization. GDM-MADDPG takes the observation of the agent as the condition of the denoising network of the diffusion model, and replaces the actor network of MADDPG with the reverse denoising process to better capture the correlation between the data, as shown in Fig. 4. In the initialization phase, each agent will establish its own Actor network and Critic network, as well as the corresponding target network, and create an experience replay buffer to store interaction data. The Actor network determines the action based on the local observation output of the agent, while the Critic network evaluates the expected reward of the state-action pairs of all agents. During training, the agent performs actions and collects environmental feedback, which is stored in the experience replay memory. In order to stabilize the training, GDM-MADDPG introduces a target network whose parameters are updated at a certain rate to achieve a smooth learning process. Next, samples are randomly drawn from the memory to update the Critic network so that it can more accurately predict the rewards of Observation-action pairs. Subsequently, the output of the Critic network is used to update the Actor network with the goal of maximizing the expected reward.
@@ 26                        Minimize the loss function to update the parameters of the Critic network
$\mathcal{A}^{n}=\{a_{1}^{n},a_{2}^{n},\ldots a_{t}^{n}\}$. In each time slot t, the agent’s action is denoted by $a_{t}^{i}$, where $i=\{1,2,...,I\}$ represents the task chosen by the agent for execution and $a_{t}^{n}=\{a_{t}^{1},a_{t}^{2},...,a_{t}^{i}\}$ means the action taken.
@@
VII Experiment Setup and Performance Evaluation
@@ VII-A Experiment Setup
We perform simulations in the following scenario: an airship positioned at an altitude of 10 km, accompanied by nine UAVs at an altitude of 3 km, covering an area of 12×12 km². The airship is located at the center of this area, whereas the leader UAV and follow UAVs are uniformly distributed throughout it. The airship obtains task directives from the ground control center, formulates the optimal flight sequence for the designated task targets, and transmits the planned route information to the lead aircraft. The lead aircraft, along with several wingmen UAV, then executes a coordinated flight along the predetermined route. It is important to note that all UAVs fall within the coverage area of the airship and are interconnected and capable of communicating with one another. Furthermore, considering overall accuracy in detecting target amounts, the Yolov5 model outperforms the seven attention-optimized Yolov5 variants, as well as other state-of-the-art object detection models [4]. In this study, we employ Yolov5 as the task model and perform experiments to evaluate the effectiveness of the proposed approach.
3) Benchmark Algorithms
@@
Figure 5: UAV path planning considering task size and flight distance.
@@ VII-B Performance Evaluation
This experiment simulated a UAV inspection area consisting of ten target coordinates, each defined by a distinct task size. Fig. 5(a) shows the proposed UAV path planning method. By comprehensively considering both task size and flight path, this method generates the global optimal path and target coordinate sequence, prioritizing target coordinates with larger task sizes while ensuring coverage of all target coordinates. In contrast, Fig. 5(b) presents a greedy algorithm which selects the best option at each step based solely on the current state, and is prone to local optima. As a result, it can lead to a significant increase in flight path length, reducing the survival time of the UAV swarm and hindering its ability to conduct long-term operations under energy constraints.
Figure 6: The training of different DNN models and benchmark algorithms.
As illustrated in Fig. 5(c), the cost associated with the optimal path finding of the UAV increases as the number of target coordinates increases. In particular, the path planning algorithm proposed in this study consistently exhibits lower costs compared to the greedy algorithm in all scenarios evaluated. Specifically, for target coordinate counts of 10, 20, 30, 40, and 50, the total cost produced by the proposed algorithm is reduced by approximately 19.3%, 21.1%, 23.0%, 25.6%, and 27.1%, respectively, compared to the greedy algorithm. This enhancement can be attributed to the expanding solution space that accompanies an increase in the number of target coordinates. As the candidate set of target coordinates grows, the proposed algorithm has greater opportunities to explore solutions that yield lower total costs, ultimately facilitating the identification of the optimal path with minimal total cost. Thus, it is evident that the algorithm developed in this study is particularly adept at addressing path planning challenges that involve a larger number of target coordinates.
@@
The convergence of GDM-MADDPG in various tasks of the DNN model is illustrated in Fig. 6(a). Despite the differences in structural complexity between DNN models, the proposed method successfully accommodates these tasks and achieves convergence. The AlexNet model, characterized by fewer layers and lower computational requirements, demonstrates favorable convergence performance. In contrast, the VGG16 model, with its greater number of layers and higher computational demands, experiences increased processing delays, resulting in lower rewards and reduced convergence efficacy for our method. Although the convergence performance of our method on the YOLOv5 model is slightly inferior to that of AlexNet, YOLOv5, as a widely adopted real-time object detection algorithm, has a model complexity that lies between the two aforementioned models. Consequently, subsequent experiments employ YOLOv5 as the experimental task model to validate the superior performance of this method in terms of AOI and task completion rate.
We perform ten independent training runs for each algorithm. As illustrated in Fig. 6(b), although the proposed GDM-MADDPG converges at a slower rate than the two algorithms, GDM-MADDPG demonstrates superior performance compared to the other solutions. The total reward attained by GDM-MADDPG increases throughout the training process, eventually converging to approximately 5000 after around 600 training episodes. This enhancement is attributed to the generative capabilities of GDM, which markedly improve action sample efficiency by progressively reducing noise through multiple denoising steps. Finally, the low performance of the two algorithms further highlights the effectiveness of the proposed GDM-MADDPG method.
Figure 7: The training of different learning rate and batch size.
By conducting a series of systematic experiments, we aim to identify the optimal values for two key parameters that influence the performance of the MADDPG algorithm: 1) learning rate and 2) batch size. The experimental results, illustrated in Fig. 7(a), clarify the convergence behavior of the reward function under varying learning rates. A learning rate of 0.01 is excessively large, preventing convergence even after 1000 training iterations. Similarly, learning rates of 0.0001 do not facilitate convergence. In particular, the 0.001 learning rate promotes faster convergence. The suitable learning rate of 0.001 is used to assess the effects of different batch sizes on training performance. As illustrated in Fig. 7(b), a batch size of 512 is recognized as the most advantageous option.
@@
This study compares the AoI of different methods in varying task sizes, as illustrated in Fig. 8(a). The AoI performance of the MADDPG-based method is significantly improved when combined with the path-planning approach. Furthermore, the GDM-MADDPG method proposed in this paper, which is based on MADDPG and incorporates the advantages of GDM in handling complex data distributions and generating high-quality task assignment decisions, improves the timeliness and generalizability of task assignment decision generation. This results in the generation of the most suitable task assignment decisions for the current state, thereby reducing the transmission time during serial task execution. According to Eq. (25), this method also addresses the optimization of path planning for target coordinates and tasks, which reduces waiting times associated with tasks and improves AoI.
The task completion rates of different methods were compared in various task sizes, as shown in Fig. 8(b). The method proposed in this paper outperforms other methods in optimizing system utility, indicating that optimized task path planning and task offloading decision generation achieve a joint optimization of multiple objectives. In terms of path planning, the method not only considers the optimal path but also accounts for the task-size requests at the target coordinates, enhancing the selectivity of local optima. Regarding DNN task assignment, factors such as task partitioning decisions and load balancing are incorporated. Additionally, the introduced GDM reverse denoising process improves the learning and generation capabilities of agent decisions, leading to well-generalized action decisions. Consequently, as the task size gradually increases, the MADDPG-based decision method lacks robust dynamic adaptability under multiple constraints. Its learning capability is constrained by the fixed learning mechanism of the actor-critic network, resulting in the absence of a strategy with generalization capabilities.
@@
The utility of different methods was compared under various task sizes, as shown in Fig. 8(c), as task sizes increase, the need for additional task assignment work is incurred under latency constraints, directly increasing the latency of task completion. This results in a downward trend in task completion rates for all methods. The proposed method considers task size constraints during the path planning stage, and task assignment decisions can be rapidly learned and generated through the GDM method, further reducing task processing delays and allowing a greater number of tasks to be completed within the specified time. In addition, these findings indicate that the proposed method is suited for large-scale tasks.
VIII Conclusion
In this paper, we addressed the challenge of planning joint flight paths, assigning DNN tasks, and balancing load in UAV networks as a two-stage optimization problem. In the first stage, we consider the task size of the area to be inspected and the shortest flight path as optimization constraints. We then employ a novel heuristic algorithm to optimize the UAV’s flight path, focusing on maximizing system utility. In the second stage, we introduce a novel GDM-MADDPG algorithm to determine optimal DNN task assignment decisions. Through numerical simulations, we demonstrate the superior performance of our proposed algorithm compared to existing benchmark solutions.